OpenStudy (plum):

How do you integrate (3^xdx)/(3^x - 7) ?

OpenStudy (amistre64):

\[\int \frac{3^x}{3^x - 7}dx \]

OpenStudy (amistre64):

the top is the derivative of the bottom except for a constant

OpenStudy (amistre64):

\[Dx(3^x)=3^x\ ln(3)\]

OpenStudy (amistre64):

multiply this by \(\cfrac{ln(3)}{ln(3)}\) and use what you need and pull out the rest as a constatnt

OpenStudy (amistre64):

\[\frac{1}{ln(3)}\ \int \frac{3^x\ ln(3)}{3^x - 7}dx=\frac{ln|3^x -7|}{ln(3)}\]

OpenStudy (amistre64):

+C ...

hero (hero):

lol...don't forget the plus see

OpenStudy (amistre64):

one way to do this is to go ... i have to use a "u" and if i dont then it cant be done ::: arms flailing and spittle flying:::

OpenStudy (anonymous):

lol spittle

OpenStudy (amistre64):

u = 3^x - 7 du = 3^x ln(3) dx dx = du/3^x ln(3) \[\int \frac{3^x}{3^x\ ln(3)\ u}du\] \[\int \frac{\cancel{3^x}}{\cancel{3^x}\ ln(3)\ u}du\] \[\int \frac{1}{\ ln(3)\ u}du\] \[\frac{1}{ln(3)}\int \frac{1}{u}du\]

OpenStudy (amistre64):

i perfer the first method, since it doesnt waste as much time :)